# Classification of unimodal isolated complete intersection singularities in positive characteristic

### v2

@@ 5.2. The classification of unimodal
If $r\geq 4,\ s\geq 6$ and $t\geq 5,$, then all the 5-jet $j_{5}(f)$ is given by $g_{ac}:=(x^{3}+axy^{4},x^{2}y+cy^{5}).$ We compute the codimension of $\widetilde{T}_{g_{ac}}(\mathcal{K}_{5}g_{ac})\cap P_{3,5}$ in $P_{3,5}$.
We show that $(y^{5},0)\notin\widetilde{T}_{g_{ac}}(\mathcal{K}_{5}g_{ac}).$ Otherwise, if $c\neq 0,\ (y^{5},0)\in\widetilde{T}_{g_{ac}}(\mathcal{K}_{5}g_{ac})$ implies $(x^{2}y,0)=e_{3}-c(y^{5},0)\in\widetilde{T}_{g_{ac}}(\mathcal{K}_{5}g_{ac}).$ Then $(x^{2}y,0)$ can only be generated by $e_{3},e_{5},e_{2},e_{4}.$ Assume $(x^{2}y,0)=re_{3}+se_{5}+t_{1}e_{2}+t_{2}e_{4}$. Then $r,s$ cannot be zero. In particular, if we write $r=r_{0}+r_{10}x+r_{01}y+r_{20}x^{2}+\dots$ and $s=s_{0}+s_{10}x+s_{01}y+s_{20}x^{2}+\dots$, we have $r_{0}\neq 0$ and $s_{01}\neq 0$. Then consider the $(0,xy^{2})$-term in $re_{3}+se_{5}+t_{1}e_{2}+t_{2}e_{4}$ and note that the coefficient must be $2s_{01}\neq 0$,contradict with $re_{3}+se_{5}+t_{1}e_{2}+t_{2}e_{4}=(x^{2}y,0).$ If $c=0$ and $a\neq 0$, $(y^{5},0)$ must be generated by $e_{3},e_{5},e_{2},e_{4}$. We can get a contradiction again by checking the $(0,xy^{2})$-term as above. If $a=c=0$, then every first component of $e_{1},\dots,e_{6}$ consists of $x$ hence $e_{1},\dots,e_{6}$ cannot generated $(y^{5},0).$
Similarly we can show $(0,y^{5})\notin\widetilde{T}_{g_{ac}}(\mathcal{K}_{5}g_{ac})$ for any $a,c\in F$. Hence $\mathrm{cod}(g_{ac})\geq 2$. By Proposition 2.19 and Proposition 2.21, $\mathcal{K}-mod(f)\geq\mathcal{K}_{5}-mod(f)\geq 2.$
@@
If $a(y)\ (resp.\ b(y),c(y))=0$, we regard $r\ (resp.\ s,t)=\infty$ here.
Now we assume $r\leq 4$ or $s\leq 6$ or $t\leq 5$ and $a(y),b(y),c(y)$ are units. We can simplify 5.11 by the trick using in [PPG24].
First we introduce recall the Implicit Function Theorem:
@@
(5.12) $F(x_{1},\dots,x_{n},0)\in\langle x_{1},\dots,x_{n}\rangle,\ \frac{\partial F}{\partial y}(x_{1},\dots,x_{n},0)\notin\langle x_{1},\dots,x_{n}\rangle,$
@@
Back to $f$ in 5.11, if
(5.13) $3r-2s\neq 0\mathrm{\ and\ }p\nmid 3r-2s,$
apply $\phi(x)=a(y)^{\frac{1}{2}}x$ and $\phi(y)=y,$ we get
@@
(5.14) $f\sim(x^{3}+xy^{r}+y^{s},x^{2}y+y^{t}\widetilde{c}(y)).$
@@
(5.16) $f\sim(x^{3}+\widetilde{a}(y)xy^{r}+y^{s},x^{2}y+y^{t}).$
@@
(5.18) $f\sim(x^{3}+xy^{r}+\widetilde{b}(y)y^{s},x^{2}y+y^{t}).$
If all 5.13, 5.15, 5.17 fail, then $r,s,t$ satisfy
(5.19) $3r-2s$ $=ap$
$3\displaystyle 2s-3t+$ 3 $=bp$
@@
But the minimal solution of 5.19 with $r\geq 3,\ s\geq 4,\ t\geq 4$ and $p\geq 5$ is exactly $r=4,\ s=6,\ t=5$, in which case $f$ is not unimodal.
@@
(i) We call the technique we use here as $\alpha,\beta$-trick, since we can easily (but not strictly) apply $\phi(x)=\alpha x,\phi(y)=\beta x$ in 5.11 and get the result. For example, when 5.13 holds, choose $\alpha,\beta$ as the solution of
(5.20) $\begin{cases}\alpha^{3}=\alpha\beta^{r}a(y)\\ \alpha\beta^{r}a(y)=\beta^{s}b(y)\end{cases}$
then apply $\phi$ on 5.11, we can get 5.14. The trick was shown in [BAM20]. The implicit function theorem provides a complete proof with a same result. But it is useless in some special characteristic, e.g. $p\mid 3r-2s$ here.
@@
[…]
@@
If $s=5$ and $r=3$, then $f\sim N_{\lambda},R_{t},P_{r,s},Z_{\lambda}$.
@@
In this case $p\nmid 3r-2s$ holds. By 5.14, we have $j_{4}(f)\sim(x^{3}+xy^{3},x^{2}y+\lambda_{0}y^{4})$, which is weighted homogeneous of degree 0 w.r.t. $(a;d)$, where $a=(3,2),d=(9,8)$. Write $g=j_{4}(f)$. Next we find $T\subset F_{a,d}^{1}R^{2}\backslash F_{a,d}^{r}R^{2}$ such that $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g)$. Then by Proposition 2.16, $f\sim g+t$ with $t\in T$.
@@
If $s=5$ and $r\geq 4$, then $f\sim N_{0}$ or $Y_{\lambda}$, where $\lambda\in\{0,1\}$.
@@
If $t=4$, let $g=j_{4}(f)=(x^{3},x^{2}y+cy^{4})\sim(x^{3},x^{2}y+y^{4})$ be the 4-jet of $f$. Then $g$ is weighted homogeneous of degree 0 w.r.t. $(a;d)$, where $a=(3,2),d=(9,8)$ as in Proposition 5.13. After the same computation, we have $F_{a,d}^{1}R^{2}\subset\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g)$. That is, $f\sim(x^{3},x^{2}y+y^{4})$.
If $t\geq 5$, let $h=(x^{3}+by^{5},x^{2}y)\sim(x^{3}+y^{5},x^{2}y)$ be the weighted 0-jet with respect to $(a,d)$, where $a=(5,3),d=(15,13)$. Computation shows $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g)$ for $T=span\langle(0,y^{5})\rangle$ (whether $p=5$ or not). Hence $f\sim(x^{3}+y^{5},x^{2}y+ay^{5})$. An obvious scaling shows $f\sim(x^{3}+y^{5},x^{2}y+\lambda y^{5})$ for $\lambda\in\{0,1\}$. ∎
@@
If $s\geq 6$ and $t=4$, then $f\sim N_{\lambda},P_{r,\infty},P_{\infty,s},P_{r,s,\lambda},\widetilde{P}_{r,s}$ or $Z_{\lambda}$.
@@
If $\lambda=12$, we write $g=j_{4}(f)=(x^{3}+12xy^{3},x^{2}y+y^{4})$, which is weighted homogeneous of degree 0 w.r.t. $(a,d)$, where $a=(3,2),d=(9,8)$. Computation shows that a weighted complete transversal is given by $(y^{5},0)$, i.e.
$F_{a,d}^{1}R^{2}\subset span\langle(y^{5},0)\rangle+\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g).$
@@
If $\lambda_{0}=1$, we write $g\sim(x^{3}+xy^{3},x^{2}y+y^{4})$. Then $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g)$ with $T=span\langle(y^{5},0),(y^{6},0),\dots,(xy^{4},0),(xy^{5},0),\dots\rangle$. Hence $f\sim(x^{3}+xy^{3}+u(y)xy^{r}+v(y)y^{s},x^{2}y+y^{4})$ with $r\geq 5,s\geq 4$.
@@
we reduce $u_{1}$ to 00. Repeat the process, we can reduce $u(y)$ to $u_{0}\in F$ and finally to $1$. Then we have $f\sim(x^{3}+xy^{3}+xy^{r},x^{2}y+y^{4})\sim P_{r,\infty}$.
@@
If $u,v\neq 0$, apply $\phi(x)=\alpha(y)^{-3},\phi(y)=\alpha(y)^{-2}$, we have $f\sim(x^{3}+xy^{3}+u(y)\alpha(y)^{2r-6}xy^{r}+v(y)\alpha(y)^{2s-9}y^{s},x^{2}y+y^{s})$. Using $\alpha,\beta$-trick, if $p\nmid 2r-2s-3$, then there exists $\alpha(y)$ such that $\alpha(y){2r-6}u(y)=\alpha(y)^{2s-9}v(y)$. By the same process to 5.21, we have $f\sim(x^{3}+xy^{3}+u_{0}xy^{r}+u_{0}y^{s},x^{2}y+y^{4})\sim(x^{3}+xy^{3}+xy^{r}+\lambda y^{s},x^{2}y+y^{4})\sim P_{r,s,\lambda}$.
@@ 6. The classification of order 3 when $charF=2$
6.1. The classification of 3-jet
Same as in Section 5, for $f=(f_{1},f_{2})$ with $\mathrm{ord}(f_{1})=3$, we have $j_{3}(f_{1})\sim(x-e_{1}y)(x-e_{2}y)(x-e_{3}y)\sim l_{1}l_{2}l_{3}$ since $F$ is algebraically closed, where $l_{i},\ i=1,2,3$ are linear forms in $R$. Here we repeat the discussion at the beginning of Section 5.
@@ 6.1. The classification of 3-jet
III.4 If $a,c\neq 0$ and $b=0$, then $j_{3}(f)\sim(x^{2}y+xy^{2},x^{3}+c^{\prime}y^{3})\sim(x^{2}y+xy^{2},x^{3}+c^{\prime\frac{1}{2}}x^{2}y+c^{\prime\frac{1}{2}}xy^{2}+c^{\prime}y^{3})\sim(x^{2}y+xy^{2},(x+cy)(x+c^{\prime\frac{1}{2}}y)^{2})$, where $c^{\prime}=\frac{c}{a}$. It’s easily to see it deduces to I or II depending on whether $c^{\prime}=1$.
@@
(6.1) $\begin{split}&(x^{3},0),\ (x^{3},x^{2}y),\ (x^{3},xy^{2}),\ (x^{3},x^{2}y+xy^{2}),\ (x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\\ &(x^{2}y,0),\ (x^{2}y,xy^{2}),\ (x^{2}y,x^{3}+xy^{2}),\ (x^{3}+x^{2}y,y^{3}+\lambda x^{2}y)\ (\lambda\neq 0),\\ &(x^{2}y,x^{3}+y^{3}),\ (x^{2}y+xy^{2},0)\end{split}$
@@ 6.2. The classification of unimodal
$\widetilde{M}_{r}^{2}$ $(x^{3}+y^{r}+ey^{l},x^{2}y+xy^{2}),\ where$ $r\geq 4,\ r\ is\ odd$
$e=e_{0}+e_{1}y^{2}+e_{2}y^{4}+\dots$ $l>r,\ l\ is\ even$
@@
(6.2) $(x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\ (x^{3}+x^{2}y,y^{3}+\lambda x^{2}y)\ (\lambda\neq 1),\ (x^{2}y,x^{3}+y^{3})$
then $j_{3}(f)$ is 3-determined. In particular, $f$ is contact equivalent to the $j_{3}(f)$ in 6.2.
@@
For $g=j_{3}(f)=(x^{2}y,x^{3}+y^{3}),\ T_{g}(\mathcal{K}g)$ is spanned by $(x^{2}y,0),(0,x^{2}y),(x^{3}+y^{3},0),(0,x^{3}+y^{3}),(0,x^{2}),(x^{2},y^{2})$, and a complete transversal is spanned by $\{(x^{r},0)\mid r\geq 4\}$. Hence $f\sim(x^{2}y+e(x)x^{r},x^{3}+y^{3})$ by Theorem 2.8 where $e(x)\in F[[x]]$ is a unit and $r\geq 4$. Applying the automorphism $\phi(x)=x,\phi(y)=y-e(x)x^{r-2},$ $f\sim(x^{2}y,x^{3}+(y+e(x)x^{r-2})^{3})\sim(x^{2}y,x^{3}+y^{3}+e(x)^{3}x^{3r-6})\sim(x^{2}y,u(x)x^{3}+y^{3}),$ where $u(x)=1+e(x)^{3}x^{3r-9}$ is a unit in $F[[x]]$. Applying $\phi(x)=x,\phi(y)=u(x)^{\frac{1}{3}}y$, then $f\sim(x^{2}y,x^{3}+y^{3})$. ∎
@@
If $j_{3}(f)\sim(x^{3},xy^{2})$, then $f\sim(x^{3}+y^{r},xy^{2})\sim M_{r},\ r\geq 4.$
@@
If $j_{3}(f)\sim(x^{3},x^{2}y+xy^{2})$, then $f\sim M_{r}^{2}$ or $\widetilde{M}_{r}^{2}$.
@@
If $r$ is even, using $\alpha,\beta$-trick we have $f\sim(x^{3}+y^{r},x^{2}y+xy^{2})$ (since $32\nmid r-$ 3).
@@
If $j_{3}(f)\sim(x^{2}y,xy^{2})$, then $f\sim(x^{2}y+y^{r},xy^{2}+y^{s})\sim L_{r,s},\ r,s\geq 4$.
@@
The complete transversal is same with Proposition 5.6, and the later process also follows from Proposition 5.6. ∎
@@
If $j_{3}(f)\sim(x^{2}y,x^{3}+xy^{2})$, then $f\sim(x^{2}y+y^{r},x^{3}+xy^{2})\sim K_{r}$, $r\geq 4$.
@@
If $i_{3}(f)\sim(x^{3},x^{2}y)$, a complete transversal is given by
$\{(xy^{r},0),(y^{s},0),(0,xy^{u}),(0,y^{v})\mid r,u\geq 3,s,v\geq 4\}$
. Then
(6.4) $f\sim(x^{3}+a(y)xy^{r}+b(y)y^{s},x^{2}y+c(y)xy^{u}+d(y)y^{v})$
with $r,u\geq 3,\ s,v\geq 4$.
@@
For $r,u\geq 4,\ s,v\geq 5$, then $j_{4}(f)=(x^{3},x^{2}y)$, and any 5-jet is of the form $g_{ac}=(x^{3}+axy^{4},x^{2}y+cxy^{3}+dy^{5})$ or $(x^{3}+axy^{4},x^{2}y+cxy^{4}+dy^{5})$ with $a,c,d\in F$. Computation or Singular program can show $(y^{5},0),(0,y^{5})\notin P_{4,5}$ in each case, hence $cod(g_{ac})=2$ for all $a,c\in F$. By Proposition 2.19 and 2.21, $\mathcal{K}-mod(f)\geq\mathcal{K}_{5}-mod(f)\geq inf\{cod(g_{ac})\}\geq 2.$
@@
If $f$ is of the form 6.4 with $s=4$, then $f$ is contact equivalent to the form $\sim(x^{3}+y^{4},x^{2}y+\mu xy^{3})$ with $\mu=0,1$ or $(x^{3}+y^{4}+xy^{3},x^{2}y+\lambda xy^{3})$ with $\lambda\in F$. That is, $f\sim X^{2}_{\mu}$ or $\widetilde{X}^{2}_{\lambda}$.
@@
In this case $f\sim(x^{3}+y^{4}+a(y)xy^{r},x^{2}y+c(y)xy^{u}+d(y)y^{v})$ after a scaler transform. Let $g=(x^{3}+y^{4},x^{2}y)$, then $g$ is weighted homogeneous of degree 0 w.r.t. $(a,d)$, where $a=(4,3),\ d=(12,11)$. Let $T=span\langle(xy^{3},0),(0,xy^{3})\rangle$, then we can check that $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}\cdot g)$, which means $f\sim(x^{3}+y^{4}+axy^{3},x^{2}y+bxy^{3})$ for $a,b\in F$. Take a scaler transform, we have $f\sim(x^{3}+y^{4},x^{2}y)$, $(x^{3}+y^{4},x^{2}y+xy^{3})$ or $(x^{3}+y^{4}+xy^{3},x^{2}y+\lambda xy^{3})$ with $\lambda\in F$. ∎

### v3

@@ II $(x^{3},y^{3}+x^{2}y)$
If $r\geq 4,\ s\geq 6$ and $t\geq 5,$ set $g=j_{4}(f)=(x^{3},x^{2}y)$, then all jets in $J_{5}^{\prime}(g)$ are equivalent to $g_{ac}:=(x^{3}+axy^{4},x^{2}y+cy^{5}),$ where $J_{5}^{\prime}(g)$ is formed by jets in $J_{5}(g)$ with $r\geq 4,\ s\geq 6$ and $t\geq 5.$
Analogously to Proposition 3.3, we can show $(y^{5},0),(0,y^{5})\notin\widetilde{T}_{g_{ac}}(\mathcal{K}_{5}g_{ac})$. Hence $\mathrm{cod}(g_{ac})\geq 2$. By Proposition 2.19 and Proposition 2.21, $\mathcal{K}-mod(f)\geq\mathcal{K}_{5}-mod(f)\geq 2.$
@@
If a(y)(resp.b(y),c(y))=0a(y)\ (resp.\ b(y),c(y))=0, we regard $r\ (resp.\ s,t)=\infty$ here.
Now we assume $r\leq 4$ or $s\leq 6$ or $t\leq 5$ and $a(y),b(y),c(y)$ are units. We can simplify (5.17) by the trick using in [PPG25].
First we recall the Implicit Function Theorem:
@@
$F(x_{1},\dots,x_{n},0)\in\langle x_{1},\dots,x_{n}\rangle,\ \frac{\partial F}{\partial y}(x_{1},\dots,x_{n},0)\notin\langle x_{1},\dots,x_{n}\rangle,$ (5.18)
@@
Back to $f$ in (5.17), if
$3r-2s\neq 0\mathrm{\ and\ }p\nmid 3r-2s,$ (5.19)
applying $\phi(x)=a(y)^{\frac{1}{2}}x$ and $\phi(y)=y,$ we get
@@
$f\sim(x^{3}+xy^{r}+y^{s},x^{2}y+y^{t}\widetilde{c}(y)),$ (5.20)
where $\widetilde{c}(y)$ is the image of $c_{2}(y)$ under the automorphism $\xi$.
@@
$f\sim(x^{3}+\widetilde{a}(y)xy^{r}+y^{s},x^{2}y+y^{t}),$ (5.22)
where $\widetilde{a}(y)$ is the image of $a(y)$ under the similar automorphism.
@@
$f\sim(x^{3}+xy^{r}+\widetilde{b}(y)y^{s},x^{2}y+y^{t}).$ (5.24)
If all (5.19), (5.21), (5.23) fail, then $r,s,t$ satisfy
$3r-2s$ $=ap,$ (5.25)
$2s-3t+3$ $=bp,$
@@
But the minimal solution of (5.25) with $r\geq 3,\ s\geq 4,\ t\geq 4$ and $p\geq 5$ is exactly r=4,s=6,t=5r=4,\ s=6,\ t=5, in which case $f$ is not unimodal.
@@
(i) We call the technique we use here as $\alpha,\beta$-trick, since we can easily (but not strictly) apply $\phi(x)=\alpha x,\phi(y)=\beta x$ in (5.17) and get the result. For example, when (5.19) holds, choose $\alpha,\beta$ as the solution of
$\begin{cases}\alpha^{3}=\alpha\beta^{r}a(y),\\ \alpha\beta^{r}a(y)=\beta^{s}b(y),\end{cases}$ (5.26)
then apply $\phi$ on (5.17), we can get (5.20). The trick was shown in [DG83]. The implicit function theorem provides a complete proof with a same result. But it is useless in some special characteristic, e.g. $p\mid 3r-2s$ here.
@@
If $s=4$, then $f\sim(x^{3}+y^{4},x^{2}y+\lambda y^{4})\sim X_{\lambda}$ in Table LABEL:table-ord3, where $\lambda=0\mathrm{\ or\ }1$.
@@
If $s=5$ and $r=3$, then $f\sim N_{\lambda},R_{t},P_{r,s},Z_{\lambda}$ in Table LABEL:table-ord3.
@@
In this case $p\nmid 3r-2s$ holds. By (5.20), we have $j_{4}(f)\sim(x^{3}+xy^{3},x^{2}y+\lambda_{0}y^{4})$, which is weighted homogeneous of degree 0 with respect to $(a;d)$, where $a=(3,2),d=(9,8)$. Write $g=j_{4}(f)$. Next we find $T\subset F_{a,d}^{1}R^{2}\backslash F_{a,d}^{r}R^{2}$ such that $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}g)$. Then by Proposition 2.16, $f\sim g+t$ with $t\in T$.
@@
If $s=5$ and $r\geq 4$, then $f\sim N_{0}$ or $Y_{\lambda}$ in Table LABEL:table-ord3, where $\lambda\in\{0,1\}$.
@@
If $t=4$, let $g=j_{4}(f)=(x^{3},x^{2}y+cy^{4})\sim(x^{3},x^{2}y+y^{4})$ be the 4-jet of $f$. Then $g$ is weighted homogeneous of degree 0 with respect to $(a;d)$, where $a=(3,2),d=(9,8)$ as in Proposition 5.13. After the same computation, we have $F_{a,d}^{1}R^{2}\subset\widetilde{T}_{g}(F^{1}\mathcal{K}g)$. That is, $f\sim(x^{3},x^{2}y+y^{4})$.
If $t\geq 5$, let $h=(x^{3}+by^{5},x^{2}y)\sim(x^{3}+y^{5},x^{2}y)$ be the weighted 0-jet with respect to $(a,d)$, where $a=(5,3),d=(15,13)$. Computation shows $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}g)$ for $T=span\langle(0,y^{5})\rangle$ (whether $p=5$ or not). Hence $f\sim(x^{3}+y^{5},x^{2}y+ay^{5})$. An obvious scaling shows $f\sim(x^{3}+y^{5},x^{2}y+\lambda y^{5})$ for $\lambda\in\{0,1\}$. ∎
@@
If $s\geq 6$ and $t=4$, then $f\sim N_{\lambda},P_{r,\infty},P_{\infty,s},P_{r,s,\lambda},\widetilde{P}_{r,s}$ or $Z_{\lambda}$ in Table LABEL:table-ord3.
@@
If $\lambda=12$, we write $g=j_{4}(f)=(x^{3}+12xy^{3},x^{2}y+y^{4})$, which is weighted homogeneous of degree 0 with respect to $(a,d)$, where $a=(3,2),d=(9,8)$. Computation shows that a weighted complete transversal is given by $(y^{5},0)$, i.e.
$F_{a,d}^{1}R^{2}\subset span\langle(y^{5},0)\rangle+\widetilde{T}_{g}(F^{1}\mathcal{K}g).$
@@
If $\lambda_{0}=1$, we write $g\sim(x^{3}+xy^{3},x^{2}y+y^{4})$. Then $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}g)$ with $T=span\langle(y^{l},0),(xy^{j},0)|l\geq 5,j\geq 4\rangle$. Hence $f\sim(x^{3}+xy^{3}+u(y)xy^{r}+v(y)y^{s},x^{2}y+y^{4})$ with $r\geq 5,s\geq 4$.
@@
we reduce $u_{1}$ to 0. Repeat the process, we can reduce $u(y)$ to $u_{0}\in F$ and finally to $1$. Then we have $f\sim(x^{3}+xy^{3}+xy^{r},x^{2}y+y^{4})\sim P_{r,\infty}$.
@@
If $u(y),v(y)\neq 0$, apply $\phi(x)=\alpha(y)^{-3},\phi(y)=\alpha(y)^{-2}$, we have $f\sim(x^{3}+xy^{3}+u(y)\alpha(y)^{2r-6}xy^{r}+v(y)\alpha(y)^{2s-9}y^{s},x^{2}y+y^{s})$. Using $\alpha,\beta$-trick, if $p\nmid 2r-2s-3$, then there exists $\alpha(y)$ such that $\alpha(y){2r-6}u(y)=\alpha(y)^{2s-9}v(y)$. By the same process to (5.27), we have $f\sim(x^{3}+xy^{3}+u_{0}xy^{r}+u_{0}y^{s},x^{2}y+y^{4})\sim(x^{3}+xy^{3}+xy^{r}+\lambda y^{s},x^{2}y+y^{4})\sim P_{r,s,\lambda}$.
@@ 6. The classification of order 3 when $charF=2$
6.1. The classification of 3-jets
Same as in Section 5, for $f=(f_{1},f_{2})$ with $\mathrm{ord}(f_{1})=3$, we have $j_{3}(f_{1})\sim(x-e_{1}y)(x-e_{2}y)(x-e_{3}y)\sim l_{1}l_{2}l_{3}$ since $F$ is algebraically closed, where $e_{1},e_{2},e_{3}$ are the roots of $j_{3}(f_{1}(x,1))$, and $l_{i},\ i=1,2,3$ are linear forms in $R$. Here we repeat the discussion at the beginning of Section 5.
@@ 6.1. The classification of 3-jets
III.4 If $a,c\neq 0$ and $b=0$, then $j_{3}(f)\sim(x^{2}y+xy^{2},x^{3}+c^{\prime}y^{3})\sim(x^{2}y+xy^{2},x^{3}+c^{\prime\frac{1}{2}}x^{2}y+c^{\prime\frac{1}{2}}xy^{2}+c^{\prime}y^{3})\sim(x^{2}y+xy^{2},(x+cy)(x+c^{\prime\frac{1}{2}}y)^{2})$, where $c^{\prime}=\frac{c}{a}$. It’s easily to see that it deduces to I or II depending on whether $c^{\prime}=1$.
@@
$\begin{split}&(x^{3},0),\ (x^{3},x^{2}y),\ (x^{3},xy^{2}),\ (x^{3},x^{2}y+xy^{2}),\ (x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\\ &(x^{2}y,0),\ (x^{2}y,xy^{2}),\ (x^{2}y,x^{3}+xy^{2}),\ (x^{3}+x^{2}y,y^{3}+\lambda x^{2}y)\ (\lambda\neq 0),\\ &(x^{2}y,x^{3}+y^{3}),\ (x^{2}y+xy^{2},0).\end{split}$ (6.1)
@@ II $(x^{3},y^{3}+x^{2}y)$
$\widetilde{M}_{r}^{2}$ $(x^{3}+y^{r}+ey^{l},x^{2}y+xy^{2}),\ where$ $r\geq 4,\ r\ is\ odd,$
$e=e_{0}+e_{1}y^{2}+e_{2}y^{4}+\dots$ l>r,lisevenl>r,\ l\ is\ even
@@
$(x^{3},y^{3}+x^{2}y),\ (x^{3},y^{3}),\ (x^{3}+x^{2}y,y^{3}+\lambda x^{2}y)\ (\lambda\neq 1),\ (x^{2}y,x^{3}+y^{3}),$ (6.2)
then $j_{3}(f)$ is 3-determined. In particular, $f$ is contact equivalent to the $j_{3}(f)$ in (6.2).
@@
For $g=j_{3}(f)=(x^{2}y,x^{3}+y^{3}),\ \widetilde{T}_{g}(\mathcal{K}g)$ is spanned by $(x^{2}y,0),(0,x^{2}y),(x^{3}+y^{3},0),(0,x^{3}+y^{3}),(0,x^{2}),(x^{2},y^{2})$}, and a complete transversal is spanned by $\{(x^{r},0)\mid r\geq 4\}$. Hence $f\sim(x^{2}y+e(x)x^{r},x^{3}+y^{3})$ by Theorem 2.8 where $e(x)\in F[[x]]$ is a unit and $r\geq 4$. Applying the automorphism $\phi(x)=x,\phi(y)=y-e(x)x^{r-2},$ $f\sim(x^{2}y,x^{3}+(y+e(x)x^{r-2})^{3})\sim(x^{2}y,x^{3}+y^{3}+e(x)^{3}x^{3r-6})\sim(x^{2}y,u(x)x^{3}+y^{3}),$ where $u(x)=1+e(x)^{3}x^{3r-9}$ is a unit in $F[[x]]$. Applying $\phi(x)=x,\phi(y)=u(x)^{\frac{1}{3}}y$, then $f\sim(x^{2}y,x^{3}+y^{3})$. ∎
@@
If $j_{3}(f)\sim(x^{3},xy^{2})$, then $f\sim(x^{3}+y^{r},xy^{2})\sim M_{r},\ r\geq 4$ in Table LABEL:table-ord3-char2.
@@
If $j_{3}(f)\sim(x^{3},x^{2}y+xy^{2})$, then $f\sim M_{r}^{2}$ or $\widetilde{M}_{r}^{2}$ in Table LABEL:table-ord3-char2.
@@
If $r$ is even, using $\alpha,\beta$-trick we have $f\sim(x^{3}+y^{r},x^{2}y+xy^{2})$ (since $2\nmid r-3$).
@@
If $j_{3}(f)\sim(x^{2}y,xy^{2})$, then $f\sim(x^{2}y+y^{r},xy^{2}+y^{s})\sim L_{r,s},\ r,s\geq 4$ in Table LABEL:table-ord3-char2.
@@
The complete transversal is the same as Proposition 5.6, and the later process also follows from Proposition 5.6. ∎
@@
If $j_{3}(f)\sim(x^{2}y,x^{3}+xy^{2})$, then $f\sim(x^{2}y+y^{r},x^{3}+xy^{2})\sim K_{r}$ in Table LABEL:table-ord3-char2, $r\geq 4$.
@@
If $j_{3}(f)\sim(x^{3},x^{2}y)$, a complete transversal is given by
$\{(xy^{r},0),(y^{s},0),(0,xy^{u}),(0,y^{v})\mid r,u\geq 3,s,v\geq 4\}.$
Then
$f\sim(x^{3}+a(y)xy^{r}+b(y)y^{s},x^{2}y+c(y)xy^{u}+d(y)y^{v}).$ (6.4)
with $r,u\geq 3,\ s,v\geq 4$
@@
For $r\geq 4,\ s\geq 6,v\geq 5$, we write $g=j_{4}(f)=(x^{3},x^{2}y)$, and any 5-jet in a open dense subset of $J^{\prime}_{5}(g)$ is of the form $g_{ac}=(x^{3}+axy^{4},x^{2}y+cxy^{3}+dy^{5})$ with $a,c,d\in F$, where $J^{\prime}_{5}(g)$ is formed by jets in $J_{5}(g)$ with $r\geq 4,s\geq 6,v\geq 5$. Computation or Singular program can show $(y^{5},0),(0,y^{5})\notin P_{4,5}$ in each case, hence $cod(g_{ac})=2$ for all $a,c\in F$. By Proposition 2.19 and 2.21, $\mathcal{K}-mod(f)\geq\mathcal{K}_{5}-mod(f)\geq inf\{cod(g_{ac})\}\geq 2.$
@@
[…]
@@
In this case $f\sim(x^{3}+y^{4}+a(y)xy^{r},x^{2}y+c(y)xy^{u}+d(y)y^{v})$ after a scaler transform. Let $g=(x^{3}+y^{4},x^{2}y)$, then $g$ is weighted homogeneous of degree 0 with respect to $(a,d)$, where a=(4,3),d=(12,11)a=(4,3),\ d=(12,11). Let $T=span\langle(xy^{3},0),(0,xy^{3})\rangle$, then we can check that $F_{a,d}^{1}R^{2}\subset T+\widetilde{T}_{g}(F^{1}\mathcal{K}g)$, which means $f\sim(x^{3}+y^{4}+axy^{3},x^{2}y+bxy^{3})$ for $a,b\in F$. Take a scaler transform, we have $f\sim(x^{3}+y^{4},x^{2}y)$, $(x^{3}+y^{4},x^{2}y+xy^{3})$ or $(x^{3}+y^{4}+xy^{3},x^{2}y+\lambda xy^{3})$ with $\lambda\in F$. ∎